Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard​ form,
The solutions are -8 and 4.

Answer by Edwin McCravy(20054)


       x = -8;    x = 4
     x+8 = 0;   x-4 = 0
       (x+8)(x-4) = 0

"FOIL" that out and finish.

Edwin